# Energy Transport Among Highly-Polarized Atoms

### v1

@@ Abstract
Harmonic ladders of energy levels are ubiquitous in nature. Beyond the paradigmatic quantum harmonic oscillator, energy ladders are the basis for many phenomena in atomic, chemical, and condensed matter systems. They emerge from vibrations in molecules and lattices [1], quantum well states in thin films [2], or as Hofstadter’s butterfly via field-induced electron motion in solids [3]. These energy structures lead to interesting studies of nonequilibrium dynamics, topological phases, and energy transport. Recent applications include memories for quantum photonic networks [4] and control of molecular disassociation through vibrational ladder climbing [5]. Manipulation of quantum confinement in quantum dots and other nanostructures can tune electronic and optical properties necessary for lasing and band gap engineering [6, 7]. Engineered higher-dimensional Hamiltonians are also possible by encoding energy levels as an additional degree of freedom, or synthetic dimension [8]. Recently, ladders of $ns$ and $np$ Rydberg levels have been mapped to a synthetic dimension to probe band structure [9].
@@
The high tunability of strong interactions among ultracold atoms make them excellent platforms for quantum simulation and studies of energy transport [10]. Recent work has explored transport in the context of quantum many-body phase transitions [11], such as the degree of correlation between entanglement growth and diffusion [12]. Energy transfer through spatial redistribution of electronic states has also been realized with direct imaging [13, 14]. On the contrary, slowed transport has been observed in the presence of quantum many-body scar states [15, 16, 17, 18], or in dynamically localized systems such as the quantum kicked rotor [19, 20, 21].
@@
In this Letter, we measure the evolution of energy transport in nearly harmonic ladders of energy levels in ultracold Rydberg atoms. The energy exchange is mediated by resonant dipole-dipole interactions among the amorphous cloud of rubidium Rydberg atoms in a magneto-optical trap (MOT). The sensitivity of Rydberg atoms to external electric and magnetic fields offers some control over the parameters of the ladder, making this a promising model system.
@@
When an atom is immersed in an electric field, the degeneracy among different angular momentum states is lifted. In non-relativistic hydrogen, the state splittings are dominated by the linear Stark effect, which leads to a harmonic spacing. While the $\ell$-degeneracy is lifted, the $m$-degeneracy is not. Thus, each rung of the ladder consists of a set of even or odd $m$-states with the number of degenerate states increasing at the center of the manifold. There are however some deviations from this in alkali atoms due to the complexity of the ion core. The quantum defects lift the $\ell$ degeneracy at zero field leading to a quadratic Stark shift at low field. Above $\ell=3$, the quantum defects are small so this quadratic shift quickly returns to a nearly linear shift as the field increases.
@@
Including fine structure and considering only the low $m_{j}$ states in rubidium, we can think about the manifold as clusters of six states each with $|m_{j}|=\frac{1}{2},\frac{3}{2},\frac{5}{2}$. For an initial cluster near the center of the manifold, dipole moments between clusters vary between 0 and 60 $ea_{0}$ whereas dipole moments within clusters can be much stronger, around $250$ $ea_{0}$. Clusters are separated by about 0.8 GHz and the initial cluster is about 14 MHz wide. The closest pair of states within the initial cluster are separated by about 0.4 MHz.
@@
Our experiment makes use of 85Rb atoms that are cooled and confined in a magneto-optical trap. With the trapping laser driving the $5s\rightarrow 5p$ cycling transition, we use two additional lasers to excite Rydberg atoms. One laser operates at 776 nm to drive the $5p\rightarrow 5d$ transition and the other excites from $5d$ to states with $nf$ character at 1265 nm.
@@
Figure 1(a) shows a Stark map for the region of interest. We must first find the desired Stark state. This is done by locking the frequency of the 1265 nm laser so that it can excite states at an energy just below the zero-field $-\ell$ $n=34$ manifold. We then measure the number of excited Rydberg atoms as the field is increased as shown in Fig. 1(b). For this experiment we chose the state that tunes into resonance at 7.15 V/cm for this excitation energy. For this initial state, neighboring clusters are nearly harmonically spaced and the spacings of states within each cluster is small. Once the desired state is found, the static electric field remains fixed at this value.
@@
Upon excitation to the Stark state, atoms can immediately begin to exchange energy through dipole-dipole interactions, which spreads population to neighboring states. To study this energy transport, we must measure the distribution of population across the Stark manifold. Selective field ionization cannot resolve the signal from these closely spaced Stark states, but it can resolve the $36d$ state from those in the $n=34$ manifold. We therefore use a pulsed microwave field as shown in Fig. 1(a) to drive population from the manifold states to the $d$-state. As the microwave frequency is scanned we measure the $d$ state population to reveal the spectrum of excited Stark states as shown in Fig. 2.
@@
Figure 2: Field ionization signal of the $36d$-state as a function of microwave frequency, shown at different delays between the end of the Rydberg excitation laser pulse and the start of the microwave pulse. Each shaded region correspond to a cluster of manifold energy levels in Fig. 1(a). Each peak within a shaded region is due to a different $36d|m_{j}|$ sublevel. All five possible peaks are most easily visible in central, red-shaded region corresponding to the initially excited state. With increasing delay, Förster interactions transfer population from the initial cluster to adjacent clusters. At a delay of 0 $\mu s$, the end of the Rydberg excitation pulse coincides with the start of the microwave pulse. This nominally indicates zero interaction time; however, the atoms interact during the 50 ns width of the excitation pulse so that transfer is already visible.
@@
Each panel in Fig. 2 shows the microwave spectrum for a particular time delay between the laser excitation pulse and the microwave probe pulse, which is the time during which atoms can interact and exchange energy. Both pulses have a width of 50 ns and a delay of 0 ns marks the time at which the trailing edge of the laser pulse is aligned with the rising edge of the microwave pulse. At 7.15 V/cm, the $36d$ state is split into five states; $j=\frac{5}{2},\ |m_{j}|=\frac{1}{2},\ \frac{3}{2},\ \frac{5}{2}$, and $j=\frac{3}{2},\ |m_{j}|=\frac{1}{2},\ \frac{3}{2}$. The range of energy between these five states is slightly smaller that the separation between neighboring manifold states. We therefore expect the microwave spectrum to show five peaks for a single Stark state cluster. The colored bands in Figs. 1 and 2 mark the range of microwave resonances associated with a single cluster of Stark states. To find the population in each Stark cluster, we integrate the signal in each band. The resulting evolution of energy transport due to dipole-dipole interaction is shown in Fig. 3 where the color and dashing of each curve matches that of Figs. 1 and 2.
Figure 3: (a) Integrated signal as a function of time for each of the highlighted manifold clusters in Figs. 1 and 2. A significant part of the asymmetry between clusters above (dashed) and below (solid) the initial cluster, visible here and in Fig. 2, is due to population fraction variation in the dipole coupling between the manifold states and the 36d states. Panel (b) uses the measured microwave coupling strengths to normalize the time dependence data.
@@
At an interaction time of $-100$ ns there should be no signal in the $d$-state since the microwave probe pulse turns off before the Rydberg excitation pulse turns on. In fact, we see signal in the $36d$-state due to the finite turnoff time of the microwave pulse. As the delay is increased to 0 ns, we see a rise in the overall signal level as well as population in neighboring Stark states due to the dipole-dipole coupling. Beyond 0 ns, the interaction continues to spread population to neighboring manifold clusters although this spread is rather slow, leaving the bulk of the population in the initial Stark state. In Fig. 3(a) we also see an asymmetry between clusters above (dashed lines) and below (solid lines) the initial state. This is largely due to the variation in the microwave coupling strength between the manifold states and the $36d$-state.
@@
To measure the microwave coupling strengths, we switched to exciting the $36d$-state [22]. We then measured, using selective field ionization, the population in the $n=34$ manifold as we scanned the microwave frequency. By separately exciting the five $36d$-states, we produced a map of the coupling strength across the manifold for each $36d$-state, which were weighted to match data shown for the initially excited state in Fig. 2. Using the same integration bins as shown in Fig. 2, we found the overall weighting factors, which were used to scale the data in Fig. 3(a). The result of this scaling, shown in Fig. 3(b), reveals a more symmetric population of clusters above and below the initially excited Stark state.
@@
Ideally, we would quench the system into resonance after excitation to the initial state. However, since the dipole-dipole interactions are resonant at such a wide range of electric fields, we cannot excite at a non-resonant field. Thus, dipole-dipole interactions occur during the 50 ns laser excitation pulse and some of the energy transport visible before 0 ns in Fig. 3 is possibly due to direct excitation of pair states [23]. However, after 0 ns the population of the initial state continues to decrease and the population of other manifold states continues to increase, a clear signature of dipole-dipole interactions. In fact, Seiler et al. determined that resonant Förster interactions among manifold states were required to model the decay of Rydberg atom population in an electrostatic trap [24].
@@
We use the calibrated density in our model of the manifold system. We numerically solve the Schrödinger equation using the two-body Hamiltonian
@@
where $\hat{\sigma}^{i}_{ab}$ is an operator that takes the $i^{th}$ atom from Stark level $a$ to Stark level $b$ and $R_{ij}$ is the distance between the $i^{th}$ and $j^{th}$ atoms. The transition dipole moments $\mu_{ab}$, which connect Stark level $a$ to $b$, are calculated using the Numerov method [26]. The first term gives the energy of each level and the second term represents represents the two-body Förster interactions. When $c=a$ and $d=b$, the second term yields hopping interactions.
In our simulations, we assume that the atoms are essentially motionless and we ignore the angular dependence of the interactions. We randomly place dozens of atoms in a spherical volume at our measured density, while using the blockade radius to estimate a minimum distance between atoms. Including enough energy levels to faithfully represent the manifold necessarily limits the number of atoms we can include in the Schrödinger equation. We iterate over each atom in the volume and choose its three nearest neighbors and then run the simulation with those four atoms so that we include the possibility of “secondary” interactions. For example, two pairs of atoms in the initial cluster could exchange energy, resulting in a pair of atoms in higher energy clusters that are now able to interact. We repeat this process for multiple volumes and average so that the final results include a few hundred simulations.
With only four atoms in the model, the simulated dynamics will be slower than in the experiment [18]. Since we also do not attempt to model direct excitation of pair states, our simulated time evolution is not suited for quantitative comparison to the experiment. However, we can look at longer times in the simulation, when even for four atoms the populations of each energy level have plateaued.
@@
As a starting point for understanding the spread of energy in our system, we consider a perfectly harmonic ladder of nine energy levels with four atoms initially in the middle level. We set the separation between energy levels to be 0.8 GHz, the mean separation between clusters in the actual experiment. The $\mu_{ab}$ are randomly drawn from a box distribution centered on the mean between-cluster dipole moments. Figure 4 shows the distribution of energy, averaged over the last five microseconds of a 100 microsecond simulation. The system has thermalized, with a relatively even energy distribution that extends over all nine energy levels. When we run this model with five or six atoms and more energy levels, we find that the system still thermalizes.
The predictions of the harmonic ladder disagree strongly with the experimental result. We can attempt to improve the model by adding some anharmonicity. Moving outward from the initial cluster used in the experiment, the mean energy of each cluster shifts by about 0.1% of the energy spacing. With this small anharmonicity, we find that the results do not significantly change and the system still thermalizes.
@@
It could be interesting to repeat this experiment for different initial clusters, or perhaps even excite individual levels within a cluster. For example, the dipole moments connecting energy levels within the cluster above our initial cluster are two orders of magnitude smaller. Exciting to this cluster would allow use to explore the importance of dipole-dipole interactions within the initial cluster.

### v2

@@ Abstract
Harmonic ladders of energy levels are ubiquitous in nature. Beyond the paradigmatic quantum harmonic oscillator, energy ladders are the basis for many phenomena in atomic, chemical, and condensed matter systems. They describe the motion of trapped ions [1, 2, 3, 4], arise from molecular and lattice vibrations [5, 6, 7], and emerge as Floquet states in strongly driven systems [8, 9, 10]. Landau levels can contribute to quantum Hall phenomena [11, 12], and have been recently observed in a variety of solid-state systems [13, 14, 15], including skyrmions [16, 17]. These energy structures lead to interesting studies of nonequilibrium dynamics, topological phases, and energy transport [18, 19, 20].
Harmonic ladders give rise to a range of modern applications, such as trapped-ion qubits for quantum information [21], or in quantum control of molecular disassociation through vibrational ladder climbing [22, 23, 24]. Manipulation of confinement in quantum dots has enabled tuning of optical and electrical properties for lasing, optoelectronic devices, and biological applications [25, 26, 27, 28, 29, 30, 31, 32, 33]. Revived interest in Floquet engineering, or coherent control of dressed electronic structures, has explored exotic quantum phases and symmetry-breaking in solid-state and atomic systems [34, 35, 36, 37, 38]. Engineered higher-dimensional Hamiltonians have been realized by encoding energy levels as an additional degree of freedom, or synthetic dimension [39]. Recently, ladders of $ns$ and $np$ Rydberg levels have been mapped to a synthetic dimension to probe band structure [20].
@@
The high tunability of strong interactions among ultracold atoms make them excellent platforms for quantum simulation and studies of thermalization. Recent work has explored energy transport in the context of quantum many-body phase transitions [40, 41] and the degree of correlation between entanglement growth and diffusion [42, 43, 44]. Energy transfer through spatial redistribution of electronic states has also been realized with direct imaging [45, 46]. Alternatively, slowed transport has been observed in the presence of quantum many-body scar states [47, 48, 49, 50], and in dynamically localized systems such as the quantum kicked rotor [51, 52, 53].
@@
In this Letter, we measure the energy transport in nearly harmonic ladders of Stark energy levels in ultracold Rydberg atoms. The energy exchange is mediated by resonant dipole-dipole interactions among the amorphous cloud of rubidium Rydberg atoms in a magneto-optical trap (MOT). The sensitivity of Rydberg atoms to external electric and magnetic fields offers some control over the parameters of the ladder, making this a promising model system.
@@
When an atom is immersed in an electric field, the degeneracy among different angular momentum states is lifted. In non-relativistic hydrogen, the state splittings are dominated by the linear Stark effect, which leads to a harmonic spacing. While the $\ell$ degeneracy is lifted, the $m$ degeneracy is not. Thus, each rung of the ladder consists of a set of even or odd $m$ states with the number of degenerate states increasing at the center of the manifold. There are however some deviations from this in alkali atoms due to the complexity of the ion core. The quantum defects lift the $\ell$ degeneracy at zero field leading to a quadratic Stark shift at low field. Above $\ell=3$, the quantum defects are small so this quadratic shift quickly returns to a nearly linear shift as the field increases.
@@
Including fine structure and considering only the low $m_{j}$ states in rubidium, we can visualize the manifold as clusters of six states each with $|m_{j}|=\frac{1}{2},\frac{3}{2},\frac{5}{2}$. For an initial cluster near the center of the manifold, dipole moments between clusters vary between 0 and 60 $ea_{0}$ whereas dipole moments within clusters can be much stronger, around $250$ $ea_{0}$. Clusters are separated by about 0.8 GHz and the initial cluster is about 14 MHz wide. The closest pair of states within the initial cluster are separated by about 0.4 MHz.
@@
Our experiment uses 85Rb atoms that are cooled and confined in a magneto-optical trap. With the trapping laser driving the $5s\rightarrow 5p$ cycling transition, we use two additional lasers to excite Rydberg atoms. One laser operates at 776 nm to drive the $5p\rightarrow 5d$ transition and the other excites from $5d$ to states with $nf$ character at 1265 nm.
@@
Figure 1(a) shows a Stark map for the region of interest. We must first find the desired Stark state. This is done by locking the frequency of the 1265 nm laser to an energy just below the zero-field $-\ell$ $n=34$ manifold. We then measure the number of excited Rydberg atoms as the field is increased as shown in Fig. 1(b). In this experiment we chose the state that tunes into resonance at 7.15 V/cm. For this initial state, neighboring clusters are nearly harmonically spaced and the spacings of states within each cluster is small. Once the desired state is found, the static electric field remains fixed at this value.
@@
Upon excitation to the Stark state, atoms can immediately begin to exchange energy through dipole-dipole interactions, which spreads population to neighboring states. To study this energy transport, we must measure the distribution of population across the Stark manifold. Selective field ionization cannot resolve the signal from these closely spaced Stark states, but it can resolve the $36d$ state from those in the $n=34$ manifold. We therefore use a pulsed microwave field as shown in Fig. 1(a) to drive population from the manifold states to the $d$ state. As the microwave frequency is scanned we measure the $d$ state population to reveal the spectrum of excited Stark states as shown in Fig. 2.
@@
Figure 2: Field ionization signal of the $36d$ state as a function of microwave frequency, shown at different delays. Each shaded region corresponds to a cluster of manifold energy levels in Fig. 1(a). Each peak within a shaded region is due to a different $36d|m_{j}|$ sublevel. All five possible peaks are most easily visible in central, red-shaded region corresponding to the initially excited state. With increasing delay, Förster interactions transfer population from the initial cluster to adjacent clusters. For 0 $\mu s$, the end of the Rydberg excitation pulse coincides with the start of the microwave pulse.
@@
Each panel in Fig. 2 shows the microwave spectrum for a particular time delay between the laser excitation pulse and the microwave probe pulse, which is the time during which atoms can interact and exchange energy. Both pulses have a width of 50 ns and a delay of 0 ns marks the time at which the trailing edge of the laser pulse is aligned with the rising edge of the microwave pulse. At 7.15 V/cm, the $36d$ state is split into five states; $j=\frac{5}{2},\ |m_{j}|=\frac{1}{2},\ \frac{3}{2},\ \frac{5}{2}$, and $j=\frac{3}{2},\ |m_{j}|=\frac{1}{2},\ \frac{3}{2}$. The range of energy between these five states is slightly smaller that the separation between neighboring manifold states. We therefore expect the microwave spectrum to show five peaks for a single Stark state cluster. The colored bands in Figs. 1 and 2 mark the range of microwave resonances associated with a single cluster of Stark states. To find the population in each Stark cluster, we integrate the signal in each band. The resulting energy transport due to dipole-dipole interaction is shown in Fig. 3 where the color and dashing of each curve matches that of Figs. 1 and 2.
[…]
@@
At an interaction time of $-100$ ns there should be no signal in the $d$ state since the microwave probe pulse turns off before the Rydberg excitation pulse turns on. In fact, we see signal in the $36d$ state due to the finite turnoff time of the microwave pulse. As the delay is increased to 0 ns, we see the overall signal rise as well as population in neighboring Stark states. Beyond 0 ns, the interaction continues to spread population to neighboring manifold clusters although this spread is rather slow, leaving the bulk of the population in the initial Stark state. In Fig. 3(a) we also see an asymmetry between clusters above (dashed lines) and below (solid lines) the initial state. This is largely due to the variation in the microwave coupling strength between manifold states and the $36d$ state.
@@
To measure the microwave coupling strengths, we switched to exciting the $36d$ state [54]. We measured, using selective field ionization, the population in the $n=34$ manifold whilst scanning microwave frequency. By separately exciting the five $36d$ states, we produced a map of the coupling strength across the manifold for each $36d$ state, which were weighted to match data shown for the initially excited state in Fig. 2. Using the same integration bins as shown in Fig. 2, we found the overall weighting factors, which were used to scale the data in Fig. 3(a). The result of this scaling, shown in Fig. 3(b), reveals a more symmetric population of clusters above and below the initially excited Stark state.
@@
Ideally, we would quench the system into resonance. However, since the dipole-dipole interactions are resonant at a wide range of electric fields, we cannot excite at a non-resonant field. Thus, interactions occur during the 50 ns laser excitation pulse and some of the energy transport visible before 0 ns in Fig. 3 is possibly due to direct excitation of pair states [55]. However, after 0 ns the initial state population continues to decrease while the populations of other clusters increase, a clear signature of dipole-dipole interactions. In fact, Seiler et al. determined that resonant Förster interactions among manifold states were required to model the decay of Rydberg atom population in an electrostatic trap [56].
@@
Using the calibrated density we model the manifold system by numerically solve the Schrödinger equation using the two-body Hamiltonian
@@
where $\hat{\sigma}^{i}_{ab}$ is an operator that takes the $i^{th}$ atom from Stark level $a$ to Stark level $b$ and $R_{ij}$ is the distance between the $i^{th}$ and $j^{th}$ atoms. The transition dipole moments $\mu_{ab}$, which connect Stark level $a$ to $b$, are calculated using the Numerov method [58]. The first term gives the energy of each level and the second term represents the two-body Förster interactions. When $c=a$ and $d=b$, the second term yields hopping interactions.
Our simulations assume that the atoms are motionless and we ignore the angular dependence of the interactions. We randomly place dozens of atoms in a spherical volume at our measured density. Including enough energy levels to faithfully represent the manifold necessarily limits the number of atoms we can include in the Schrödinger equation. We iterate over each atom in the volume and choose its three nearest neighbors and then run the simulation with those four atoms so that we include the possibility of “secondary” interactions. For example, two pairs of atoms in the initial cluster could exchange energy, resulting in a pair of atoms in other energy clusters that are now able to interact. We repeat this process for multiple volumes and average so that the final results include a few hundred simulations.
With only four atoms in the model, the simulated dynamics will be slower than in the experiment [50]. Since we also do not model direct excitation of pair states, our simulated evolution is not suited for quantitative comparison to the experiment. However, we can look at longer simulated times, when the populations of each energy cluster have plateaued.
@@
As a starting point for understanding the spread of energy in our system, we consider a perfectly harmonic ladder of nine energy levels with four atoms initially in the middle level. We set the separation between energy levels to be 0.8 GHz, the mean separation between clusters in the actual experiment. The $\mu_{ab}$ are randomly drawn from a box distribution centered on the mean between-cluster dipole moments. Figure 4 shows the distribution of energy, averaged over the last five microseconds of a 100-microsecond simulation. The system has thermalized, with a relatively even energy distribution. When we run this model with five or six atoms and more energy levels, we find that the system still thermalizes.
The predictions of the harmonic ladder disagree strongly with the experimental result. We can attempt to improve the model by adding some anharmonicity. Moving outward from the initial cluster used in the experiment, the mean energy of each cluster shifts by about 0.1% of the cluster spacing. With this small anharmonicity, we find that the results do not significantly change and the system still thermalizes.